OpenStudy (anonymous):

find x- and y- intercepts and graph the equation plotting the intercepts. -2x + 7y = -7

OpenStudy (anonymous):

first arrange the equation according to y = mx+b -2x + 7y = -7 7y = 2x - 7 divide both sides by 7 y = (2/7)x - 1 now for solving for x-intercept, substitute 0 to y. and for solving for y-intercept, substitute 0 to x.